If the saving rate is 1 (i.e., s = 1), we know that:

a. Capital per worker, K/N = 0.

b. Output per worker, Y/N = depreciation per worker.

c. Consumption per worker, C/N, will be at its highest level.

d. Output per worker, Y/N = 0.
Answer #1

Answer

The correct answer is (b) Output per worker, Y/N = depreciation per worker.

Formula :

where = Change in capital per worker, s = saving rate = 1, y = Y/N = Output/worker and d = Depreciation per worker

Steady state occurs when = 0

Hence we have : = 0

=> sy = d

As s = 1 => 1*y = d

=> y = D

=> Output per per worker = Depreciation per worker.

Thus, When s = saving rate = 1 and steady state occurs then Output per per worker = Depreciation per worker.

Hence, the correct answer is (b) Output per worker, Y/N = depreciation per worker.